# Topology-induced symmetry breaking demonstrated in antiferromagnetic magnons on a Möbius strip

### v4

@@ Abstract
We propose a mechanism of topology-induced symmetry breaking, where certain local symmetry preserved by the Hamiltonian is explicitly broken in the eigenmodes of excitation due to topologically nontrivial boundary conditions. We demonstrate this phenomenon by studying magnonic excitations on a Möbius strip comprising of two antiferromagnetically coupled spin chains. Even with a simple Hamiltonian respecting local rotational symmetry and all local curvature effects ignored, magnons exhibit linear polarization of the Néel vector devoid of chirality, forming two non-degenerate branches that cannot be smoothly connected to or be decomposed by the circularly-polarized magnons. Correspondingly, one branch undergoes a spectral shift and only admits standing waves of half-integer wavelength, whereas the other only affords standing waves of integer wavelength. The suppression of chirality in the order parameter dynamics, hence the pattern of standing waves, can be generalized to other elementary excitations on non-orientable surfaces. Under the Möbius boundary condition, we further identify an exotic phase hosting spontaneous antiferromagnetic order whilst all exchange couplings are ferromagnetic. Our findings showcase the profound influence of real-space topology on the physical nature of not just the ground state but also the quasiparticles.
@@ II Model and Solutions
Consider a nano-ribbon with the AFM configuration shown in Fig. 2(a), where the red and blue arrows indicate the spins on the A and B sub-lattices in the ground state. A Möbius strip can then be constructed based on either Fig. 2(b) or Fig. 2(c), which are topologically distinct and satisfy different TBCs to be specified later. Excluding the geometric effects originating from the curvature of the strip [23, 24, 25, 26, 27, 28, 29], we adopt a minimal Hamiltonian that preserves the local rotational symmetry about the local $z$ axis:
@@
where $i$ labels the lattice on the strip, $J_{F}$ ($J_{AF}$) is the nearest-neighbor exchange interaction for (between) the same (different) spin species, and $K$ is the perpendicular easy-axis anisotropy. The collinear AFM ordering depicted in Fig. 2 is well protected by the magnetic anisotropy $K$, which suppresses quantum fluctuations in the reduced dimension. The summation $\langle i,j\rangle$ includes the link between 1 and $N$ such that everywhere on the strip looks the same regardless of the numbering. In our convention, all parameters are positive. Contrary to a single ferromagnetic spin chain arranged on a Möbius strip [23, 24], our system is free from geometrical frustration, thus no domain wall is present.
@@
where $x_{i}$ specifies the position of site $i$ from $1$ to $N$ along the strip [see Fig. 2], $L$ is the total length of the nano-ribbon, and $\xi=\pm 1$ corresponds to the two distinct ways of connection illustrated in Fig. 2(b) and 2(c). The new magnon operators $\alpha_{i}$ and $\beta_{i}$ satisfy not only the bosonic commutation relations but also the PBCs
@@
$\mathcal{H}_{\alpha(\beta)}=S\begin{pmatrix}Q_{\alpha(\beta)}&-J_{AF}/2\\ -J_{AF}/2&Q_{\alpha(\beta)}\end{pmatrix},$ (14)
where $Q_{\alpha}=K+J_{AF}/2+J_{F}[1-\cos{(k-\xi\pi/L)l}]$ and $Q_{\beta}=K+J_{AF}/2+J_{F}[1-\cos{(kl)}]$. It should be noted that in the $\alpha$ sector, magnons of momentum $k$ couple those of momentum $-k+2\pi\xi/L$; whereas in the $\beta$ sector, $k$ pairs with $-k$ without a shift.
Owing to the bosonic commutation relations of the BdG basis, we need to diagonalize $\sigma_{z}\mathcal{H}_{\alpha(\beta)}$ rather than $\mathcal{H}_{\alpha(\beta)}$ for the magnon solutions [30, 31]. In this regard, we obtain the eigen-frequencies (we set $\hbar=1$)
@@
$v_{\alpha(\beta)}^{\pm}=$ $\left(\sqrt{q_{\alpha(\beta)}^{1}}\pm\sqrt{q_{\alpha(\beta)}^{2}},\sqrt{q_{\alpha(\beta)}^{1}}\mp\sqrt{q_{\alpha(\beta)}^{2}}\right)^{\mathrm{T}},$ (16)
@@
$\tilde{\alpha}_{k}$ $=\frac{\sqrt{q_{\alpha}^{1}}+\sqrt{q_{\alpha}^{2}}}{2\sqrt{Q_{\alpha}}}\alpha_{k}+\frac{\sqrt{q_{\alpha}^{1}}-\sqrt{q_{\alpha}^{2}}}{2\sqrt{Q_{\alpha}}}\alpha_{-k+2\pi/L}^{\dagger}$ (17a)
$\tilde{\beta}_{k}$ $=\frac{\sqrt{q_{\beta}^{1}}+\sqrt{q_{\beta}^{2}}}{2\sqrt{Q_{\beta}}}\beta_{k}+\frac{\sqrt{q_{\beta}^{1}}-\sqrt{q_{\beta}^{2}}}{2\sqrt{Q_{\beta}}}\beta_{-k}^{\dagger}$ (17b)
and their $\tilde{\alpha}_{k}^{\dagger}$, $\tilde{\beta}_{k}^{\dagger}$ counterparts. Figure 3(a) and (b) plot the discretized dispersion relations for $N=10$ (only the lowest few states on each branch are shown), along with illustrations of the magnon eigenmodes at $k=0$. While the $\beta$ modes distribute symmetrically $\omega_{\beta}(-k)=\omega_{\beta}(k)$, the $\alpha$ branch shifts rightward by $\delta k=\pi/L$ such that $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$. The skewed $\omega_{\alpha}(k)$ is intimately related to the asymmetric paring of $\Psi_{\alpha}$ in Eq. (13), which originates from the non-trivial topology of the Möbius strip. It is easy to verify that setting $\xi=-1$ [i.e., using the connection of Fig. 2(c)] leads to a leftward shift of $\omega_{\alpha}(k)$, or $\delta k=-\pi/L$. Interestingly, if we reversely count the sites on the strip, the spectral shift $\delta k$ also flips sign, but in this case the eigenvectors are different from what one would obtain for $\xi=-1$.
@@
$\tilde{\alpha}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}(k-\delta k)x_{i}}}{2\sqrt{NSQ_{\alpha}}}\left[\sqrt{q_{\alpha}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\alpha}^{2}}S_{Ai}^{y}\right.$
@@
$\tilde{\beta}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}kx_{i}}}{2\sqrt{NSQ_{\beta}}}\left[\sqrt{q_{\beta}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Ai}^{y}\right.$
$\qquad\qquad\qquad\ \left.+\sqrt{q_{\beta}^{1}}S_{Bi}^{x}+\mathrm{i}\sqrt{q_{\beta}^{2}}S_{Bi}^{y}\right].$ (18b)
@@
$\bm{S}_{A/B}^{\beta}$ $\sim\mathrm{Re}\!\left[\left(\sqrt{q_{\beta}^{1}}\hat{x}\pm\mathrm{i}\sqrt{q_{\beta}^{2}}\hat{y}\right)e^{\mathrm{i}(\omega_{\beta}t-kx)}\right],$ (19b)
@@
for the $\alpha$- and $\beta$-branch, respectively, where the $+$ (--) sign corresponds to the A (B) sublattice. Because $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$, both the $\alpha$ and $\beta$ branches exhibit left-handed (right-handed) elliptical precession of $\bm{S}_{A}$ ($\bm{S}_{B}$) with the major axes of the trajectory lying along $x$; $\bm{S}_{A}$ and $\bm{S}_{B}$ always precess about the easy-axis with the same amplitude but opposite chirality. As a result, the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ exhibits linear oscillation devoid of chirality, on which we will further elaborate in the next section.
@@
Figure 3: (a, b) Illustrations of the $\alpha$ and $\beta$ modes at $k=0$, and the plots of the lowest few modes on each branch for $\xi=1$, $N=10$, $l=1$, $S=2$, $J_{F}=J_{AF}=1$, and $K=0.1$. While $\omega_{\beta}(-k)=\omega_{\beta}(k)$ is symmetric, $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$ is skewed by $\delta k=\pi/L$. The spectral shift of the $\alpha$ branch is accompanied by an intrinsic $\pi$ phase difference in the spin precessions between sites $i=1$ and $i=N$, reflecting the impact of the Möbius TBC. (c) Left: a 3D illustration of the spin precessions in the $\alpha$ and $\beta$ modes, where $\bm{S}_{A}$ and $\bm{S}_{B}$ rotate elliptically with an equal amplitude and opposite chirality, rendering the Néel vector linearly-polarized (see the color-shaded planes). Right: an illustration of the oscillating Néel vector (solid blue) and the oscillating spin vector (dashed red) for the $\alpha$ and $\beta$ modes based on Eq. (20). That $\delta\bm{n}_{\alpha}\perp\delta\bm{n}_{\beta}$ should not be confused by their being shown in different locations. (d) Ordinary right-circular ($R$) and left-circular ($L$) magnon modes, and their superposition $L-R$ and $L+R$ characterized by a linearly-polarized Néel vector (solid blue) orthogonal to the oscillating spin vector (dashed red).
@@
The distinctions between the two branches manifest in two aspects, as demonstrated in Fig. 3(a)–(b). First, from a local bird-eye view, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\alpha$-mode at $k=0$ overlap with each other when passing the minor axes of their elliptical trajectories while becoming back-to-back when passing the major axes. On the contrary, $\bm{S}_{A}$ and $\bm{S}_{B}$ in the $\beta$-mode at $k=0$ are back-to-back on the minor axes while overlapping on the major axes. Second, due to the momentum shift $\delta k=\pi/L$ in the $\alpha$ branch, the spin precessions on site $i=N$ differ from those on $i=1$ by a $\pi$ phase even for $k=0$, which exactly compensates the impact of the Möbius TBC that connects sites $1$ and $N$ with a half twist. In contrast, the $\beta$ mode at $k=0$ does not exhibit a phase difference between $1$ and $N$, which, in combination with the first distinction, is just commensurate with the TBC.
@@ III Analysis and Discussion
To further understand the unique characteristics of the magnon eigenmodes obtained above, we draw a 3D perspective in Fig. 3(c) where $\bm{S}_{A}$ and $\bm{S}_{B}$ share the same origin such that their precessional trajectories are concentric about the local $z$ axis. It is easy to deduce that the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ undergoes a pendulum-like oscillation confined on the plane containing the major (minor) axes of the two elliptical trajectories in an $\alpha$ mode ($\beta$ mode). Comparatively, the total spin vector $\bm{S}=(\bm{S}_{A}+\bm{S}_{B})/2S$ oscillates linearly on a plane orthogonal to that of $\bm{n}$. Per the definition of spin wave polarization (in terms of the order parameter dynamics) [19, 20, 21], both the $\alpha$ and $\beta$ modes are linearly polarized, thus being devoid of chirality.
@@
$\delta\bm{n}_{\beta}(k)$ $\sim\hat{y}\sqrt{q_{\beta}^{2}}\sin{[\omega_{\beta}(k)t-kx]},$ (20c)
$\bm{S}_{\beta}(k)$ $\sim\hat{x}\sqrt{q_{\beta}^{1}}\cos{[\omega_{\beta}(k)t-kx]},$ (20d)
all of which are indeed linearly polarized bearing null chirality, confirming the picture inferred in Fig. 3(c). As $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$, the oscillation amplitude of the Néel vector is larger (smaller) than that of the total spin in an $\alpha$ ($\beta$) mode, namely, $|\delta\bm{n}_{\alpha}|>|\bm{S}_{\alpha}|$ and $|\delta\bm{n}_{\beta}|<|\bm{S}_{\beta}|$. The geometry embedded in Eq. (20) is illustrated by the right panel of Fig. 3(c).
@@
where $n=1,3,5...$ is odd integer representing the number of nodes as shown in Fig. 4(a). For an even integer, it is impossible to separate $t$ and $x$ because of the spectral shift. On the other hand, mode pairing within the $\beta$ branch can only happen between $k_{n\pm}^{\beta}=\pm n\delta k$ with an even integer $n=0,2,4...$ such that
@@
At this point, it is instructive to compare the unique magnon eigenmodes on a Möbius strip with what would become the eigenmodes if the topologically trivial PBC is imposed on the nano-ribbon in Fig. 2(a). Under the PBC, the nano-ribbon only admits the well-known eigenmodes in collinear AFM materials [19, 20, 21] since we have excluded the local curvature effect from the Hamiltonian. As illustrated in Fig. 3(d), these modes are the right-circular ($R$) and left-circular ($L$) eignmodes, which, by a coherent superposition, can form $L-R$ and $L+R$ featuring elliptical precessions of $\bm{S}_{A}$ and $\bm{S}_{B}$ with opposite chirality, leading to a linearly-polarized oscillation of the Néel vector $\bm{n}$ with null chirality (so is $\bm{S}$). While $L-R$ and $L+R$ are linearly polarized, the eigenspace they span is topologically distinct from that spanned by the $\alpha$ and $\beta$ modes. To be specific, while $L-R$ is locally identical to the $\alpha$ mode, it is not accompanied by a built-in $\pi$ phase shift at $k=0$, let alone a spectral shift. The $\beta$ modes do not even locally resemble $L-R$ or $L+R$; they are emergent eigenmodes engendered by the Möbius topology. Unlike $L-R$ and $L+R$, neither $\alpha$ nor $\beta$ can be expressed as a linear superposition of $R$ and $L$. This is because the Möbius strip is a non-orientable manifold on which the chirality of spin precession becomes ambiguous globally, given that $+z$ and $-z$ are indistinguishable. In other words, the eigenspace spanned by $\alpha$ and $\beta$ is not smoothly connected to that spanned by $R$ and $L$; the two cases feature distinct topological sectors.
The structure of the eigenspace deserves an intuitive explanation. The spin precession on each sublattice can be decomposed into $\ket{\circlearrowright}$ and $\ket{\circlearrowleft}$, while the sublattice degree of freedom is represented by |$\ket{A}$ and |$\ket{B}$. With all constraints relieved, the linear space spanned by the direct products of these vectors is 4D. Under the impact of TBC, however, the magnon eigenspace only covers a 2D subspace. Specifically, the quanta of $\alpha$ and $\beta$ eigenmodes described by Eqs. (17) and (18) can be recast into a suggestive from
@@
where |$\absolutevalue{u_{\rm small}^{\alpha(\beta)}}<\absolutevalue{u_{\rm large}^{\alpha(\beta)}}$ | with $u_{\rm large}^{\alpha(\beta)}+u_{\rm small}^{\alpha(\beta)}\propto\sqrt{q_{\alpha(\beta)}^{1}}$ and $u_{\rm large}^{\alpha(\beta)}-u_{\rm small}^{\alpha(\beta)}\propto\sqrt{q_{\alpha(\beta)}^{2}}$. Comparatively, the eigenspace under the constraint of ordinary PBC, as illustrated in Fig. 3(d), is spanned by

### v5

@@ Abstract
We propose a mechanism of topology-induced symmetry breaking, where certain local symmetry preserved by the Hamiltonian is explicitly broken in the eigenmodes of excitation due to nontrivial real-space topology. We demonstrate this phenomenon by studying magnonic excitations on a Möbius strip comprising of two antiferromagnetically coupled spin chains. Even with a simple Hamiltonian respecting local rotational symmetry and all local curvature effects ignored, magnons exhibit linear polarization of the Néel vector devoid of chirality, forming two non-degenerate branches that cannot be smoothly connected to or be globally decomposed by the circularly-polarized magnons. Correspondingly, one branch undergoes a spectral shift and only admits standing waves of half-integer wavelength, whereas the other only affords standing waves of integer wavelength. Under the Möbius boundary condition, we further identify an exotic phase hosting spontaneous antiferromagnetic order whilst all exchange couplings are ferromagnetic. The suppression of chirality in the order parameter dynamics, hence the pattern of standing waves, can be generalized to other elementary excitations on non-orientable surfaces. Our findings showcase the profound influence of real-space topology on the physical nature of not just the ground state but also the quasiparticles.
@@ II Model and Solutions
Consider a nano-ribbon with the AFM configuration shown in Fig. 2(a), where the red and blue arrows indicate the spins on the A and B sub-lattices in the ground state. A Möbius strip can then be constructed based on either Fig. 2(b) or Fig. 2(c), which are topologically distinct and satisfy different TBCs to be specified later. To unambiguously separate the influence of real-space topology from local geometric effects, we intentionally exclude the curvature of the strip [23, 24, 25, 26, 27, 28, 29]. In this regard, we adopt a minimal Hamiltonian that preserves the local rotational symmetry about the local $z$ axis:
@@
where $i$ labels the lattice on the strip, $J_{F}$ ($J_{AF}$) is the nearest-neighbor exchange interaction for (between) the same (different) spin species, and $K$ is the perpendicular easy-axis anisotropy. In our convention, all parameters are positive. The collinear AFM ordering depicted in Fig. 2 is well protected by the magnetic anisotropy $K$, which suppresses quantum fluctuations in the reduced dimension. The summation $\langle i,j\rangle$ is equivalent on all links, including that between $1$ and $N$, such that every link in the longitudinal direction looks the same; we are free to re-number the whole strip by moving $1$ to anywhere. Contrary to a single ferromagnetic spin chain arranged on a Möbius strip [23, 24], our system is free from geometrical frustration, thus no domain wall is present.
@@
where $x_{i}$ specifies the position of site $i$ along the strip [see Fig. 2], $L$ is the total length of the nano-ribbon, and $\xi=\pm 1$ corresponds to the two distinct ways of connection illustrated in Fig. 2(b) and 2(c). Now, the new operators $\alpha_{i}$ and $\beta_{i}$ satisfy not only the bosonic commutation relations but also the PBCs
@@
$\mathcal{H}_{\alpha(\beta)}=S\begin{pmatrix}Q_{\alpha(\beta)}&\mp J_{AF}/2\\ \mp J_{AF}/2&Q_{\alpha(\beta)}\end{pmatrix},$ (14)
where $Q_{\alpha}=K+J_{AF}/2+J_{F}[1-\cos{(k-\xi\pi/L)l}]$ and $Q_{\beta}=K+J_{AF}/2+J_{F}[1-\cos{(kl)}]$ and the $+$ (--) sign in the off-diagonal elements corresponds to the $\mathcal{H}_{\alpha}$ ($\mathcal{H}_{\beta}$). It should be noted that in the $\alpha$ sector, magnons of momentum $k$ couple those of momentum $-k+2\pi\xi/L$; whereas in the $\beta$ sector, $k$ pairs with $-k$ without a shift.
Owing to the bosonic commutation relations of the BdG basis, we need to diagonalize $\sigma_{z}\mathcal{H}_{\alpha(\beta)}$ rather than $\mathcal{H}_{\alpha(\beta)}$ for the magnon solutions [30]. This can properly take care of the “minus sign” associated with the Bogoliubov normalization for bosons [31]. By doing so, we obtain the eigen-frequencies (we set $\hbar=1$)
@@
$v_{\alpha}^{\pm}=$ $\left(\sqrt{q_{\alpha}^{1}}\pm\sqrt{q_{\alpha}^{2}},\sqrt{q_{\alpha}^{1}}\mp\sqrt{q_{\alpha}^{2}}\right)^{\mathrm{T}},$ (16a)
$v_{\beta}^{\pm}=$ $\left(\sqrt{q_{\beta}^{1}}\pm\sqrt{q_{\beta}^{2}},-\sqrt{q_{\beta}^{1}}\pm\sqrt{q_{\beta}^{2}}\right)^{\mathrm{T}},$ (16b)
@@
$\tilde{\alpha}_{k}$ $=\frac{\sqrt{q_{\alpha}^{1}}+\sqrt{q_{\alpha}^{2}}}{2(q_{\alpha}^{1}q_{\alpha}^{2})^{\frac{1}{4}}}\alpha_{k}+\frac{\sqrt{q_{\alpha}^{1}}-\sqrt{q_{\alpha}^{2}}}{2(q_{\alpha}^{1}q_{\alpha}^{2})^{\frac{1}{4}}}\alpha_{-k+2\pi/L}^{\dagger}$ (17a)
$\tilde{\beta}_{k}$ $=\frac{\sqrt{q_{\beta}^{1}}+\sqrt{q_{\beta}^{2}}}{2(q_{\beta}^{1}q_{\beta}^{2})^{\frac{1}{4}}}\beta_{k}-\frac{\sqrt{q_{\beta}^{1}}-\sqrt{q_{\beta}^{2}}}{2(q_{\beta}^{1}q_{\beta}^{2})^{\frac{1}{4}}}\beta_{-k}^{\dagger}$ (17b)
and their counterparts $\tilde{\alpha}_{k}^{\dagger}$, $\tilde{\beta}_{k}^{\dagger}$. Figure 3(a) and (b) plot the discretized dispersion relations for $N=10$ (only the lowest few states on each branch are shown), along with illustrations of the magnon eigenmodes at $k=0$. While the $\beta$ modes distribute symmetrically $\omega_{\beta}(-k)=\omega_{\beta}(k)$, the $\alpha$ branch shifts rightward by $\delta k=\pi/L$ such that $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$. The skewed $\omega_{\alpha}(k)$ is intimately related to the asymmetric paring of $\Psi_{\alpha}$ in Eq. (13), which originates from the non-trivial topology of the Möbius strip. It is easy to verify that setting $\xi=-1$ [i.e., using the connection of Fig. 2(c)] leads to a leftward shift of $\omega_{\alpha}(k)$, or $\delta k=-\pi/L$. Interestingly, if we reversely count the sites on the strip, the spectral shift $\delta k$ also flips sign, but in this case the eigenvectors are different from what one would obtain for $\xi=-1$.
@@
$\tilde{\alpha}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}(k-\delta k)x_{i}}}{2(q_{\alpha}^{1}q_{\alpha}^{2})^{\frac{1}{4}}\sqrt{NS}}\left[\sqrt{q_{\alpha}^{1}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\alpha}^{2}}S_{Ai}^{y}\right.$
@@
$\tilde{\beta}_{k}^{\dagger}=$ $\sum_{i}\frac{e^{\mathrm{i}kx_{i}}}{2(q_{\beta}^{1}q_{\beta}^{2})^{\frac{1}{4}}\sqrt{NS}}\left[\sqrt{q_{\beta}^{2}}S_{Ai}^{x}-\mathrm{i}\sqrt{q_{\beta}^{1}}S_{Ai}^{y}\right.$
$\qquad\qquad\qquad\ \left.+\sqrt{q_{\beta}^{2}}S_{Bi}^{x}+\mathrm{i}\sqrt{q_{\beta}^{1}}S_{Bi}^{y}\right].$ (18b)
@@
$\bm{S}_{A/B}^{\beta}$ $\sim\mathrm{Re}\!\left[\left(\sqrt{q_{\beta}^{2}}\hat{x}\pm\mathrm{i}\sqrt{q_{\beta}^{1}}\hat{y}\right)e^{\mathrm{i}(\omega_{\beta}t-kx)}\right],$ (19b)
@@
for the $\alpha$- and $\beta$-branch, respectively, where the $+$ (--) sign corresponds to the A (B) sublattice. Because $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$, both the $\alpha$ and $\beta$ branches exhibit left-handed (right-handed) elliptical precession of $\bm{S}_{A}$ ($\bm{S}_{B}$). The major axes of the trajectory for the $\alpha$ ($\beta$) mode lies along $x$ ($y$); $\bm{S}_{A}$ and $\bm{S}_{B}$ always precess about the easy-axis with the same amplitude but opposite chirality. As a result, the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ exhibits linear oscillation devoid of chirality, on which we will further elaborate in the next section.
@@
Figure 3: (a, b) Illustrations of the $\alpha$ and $\beta$ modes at $k=0$, and the plots of the lowest few modes on each branch for $\xi=1$, $N=10$, $l=1$, $S=2$, $J_{F}=J_{AF}=1$, and $K=0.1$. While $\omega_{\beta}(-k)=\omega_{\beta}(k)$ is symmetric, $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$ is skewed by $\delta k=\pi/L$. The spectral shift of the $\alpha$ branch is accompanied by an accumulative $\pi$ phase in the spin precessions as we move from site $i$ to $i+N$ (i.e., traveling by a full loop), reflecting the impact of the Möbius TBC. (c) Left: a 3D illustration of the spin precessions in the $\alpha$ and $\beta$ modes, where $\bm{S}_{A}$ and $\bm{S}_{B}$ rotate elliptically with an equal amplitude and opposite chirality, rendering the Néel vector linearly-polarized. Right: illustration of the oscillating Néel vector (solid blue) and the oscillating spin vector (dashed red) for the $\alpha$ and $\beta$ modes based on Eq. (20). That $\delta\bm{n}_{\alpha}\perp\delta\bm{n}_{\beta}$ should not be confused by their being shown in different locations. (d) Ordinary AFM right-circular ($R$) and left-circular ($L$) magnon modes, and their superposition $L+R$ and $L-R$ characterized by a linearly-polarized Néel vector (solid blue) orthogonal to the oscillating spin vector (dashed red).
@@
As demonstrated in Fig. 3(a)–(b), the two branches are quite different in character even though they share one thing in common: from a local bird-eye view, $\bm{S}_{A}$ and $\bm{S}_{B}$ in both $\alpha$ and $\beta$ mode at $k=0$ overlap with each other when passing the minor axes of their elliptical trajectories while becoming back-to-back when passing the major axes. Due to the momentum shift $\delta k=\pi/L$ in the $\alpha$ branch, the spin precessions on site $i=N+1$ (same as $i=1$) acquires a $\pi$ phase (accumulated through every site) even for $k=0$, which exactly compensates the impact of the Möbius TBC that connects sites $1$ and $N$ with a half twist. In contrast, the $\beta$ mode at $k=0$ does not exhibit any phase difference between $1$ and $N$, which is just commensurate with the TBC. From a local perspective (disregarding the phase difference among different sites), the $\alpha$ and $\beta$ modes have their major axes of spin precessions perpendicular to each other, rendering the planes of Néel vector oscillation orthogonal.
@@ III Analysis and Discussion
To further understand the unique characteristics of the magnon eigenmodes obtained above, we draw a 3D perspective in Fig. 3(c) where $\bm{S}_{A}$ and $\bm{S}_{B}$ share the same origin such that their precessional trajectories are concentric about the local $z$ axis. It is easy to deduce that the Néel vector $\bm{n}=(\bm{S}_{A}-\bm{S}_{B})/2S$ undergoes a pendulum-like oscillation confined on the plane containing the major axes of the two elliptical trajectories. Comparatively, the total spin vector $\bm{S}=(\bm{S}_{A}+\bm{S}_{B})/2S$ oscillates linearly on a plane orthogonal to that of $\bm{n}$. Per the definition of spin wave polarization (in terms of the order parameter dynamics) [19, 20, 21], both the $\alpha$ and $\beta$ modes are considered linearly polarized, thus being devoid of chirality.
@@
$\delta\bm{n}_{\beta}(k)$ $\sim\hat{y}\sqrt{q_{\beta}^{1}}\sin{[\omega_{\beta}(k)t-kx]},$ (20c)
$\bm{S}_{\beta}(k)$ $\sim\hat{x}\sqrt{q_{\beta}^{2}}\cos{[\omega_{\beta}(k)t-kx]},$ (20d)
all of which are indeed linearly polarized bearing null chirality, confirming the picture inferred in Fig. 3(c). As $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}>0$, the oscillation amplitude of the Néel vectors are larger than that of the total spin in both modes, namely, $|\delta\bm{n}_{\alpha(\beta)}|>|\bm{S}_{\alpha(\beta)}|$. The geometrical relations embedded in Eq. (20) are illustrated by the right panel of Fig. 3(c).
@@
where $n=1,3,5...$ is odd integer representing the number of nodes as shown in Fig. 4(a). For an even integer, it is impossible to separate $t$ and $x$ because of the spectral shift and its ensuing $\pi$-phase shift in the spin precessions. On the other hand, mode pairing within the $\beta$ branch can only happen between $k_{n\pm}^{\beta}=\pm n\delta k$ with an even integer $n=0,2,4...$ such that
@@
At this point, it is instructive to compare the unique magnon eigenmodes on a Möbius strip with what would become the eigenmodes if the topologically trivial PBC is imposed on the nano-ribbon in Fig. 2(a). Under the PBC, the nano-ribbon only admits the well-known eigenmodes in collinear AFM materials [19, 20, 21] since we have excluded the local curvature effect from the Hamiltonian. As illustrated in Fig. 3(d), these modes are the right-circular ($R$) and left-circular ($L$) eignmodes, which, by a coherent superposition, can form $L+R$ and $L-R$ featuring elliptical precessions of $\bm{S}_{A}$ and $\bm{S}_{B}$ with opposite chirality, leading to a linearly-polarized oscillation of the Néel vector $\bm{n}$ with null chirality (so is $\bm{S}$). While $L+R$ and $L-R$ are linearly polarized, the eigenspace they span is topologically distinct from that spanned by the $\alpha$ and $\beta$ modes. To be specific, while $L+R$ is locally identical to the $\alpha$ mode, it is not accompanied by a built-in $\pi/L$ phase shift at $k=0$ for every location, let alone a spectral shift. Despite their superficial similarity to $(L-R)$ and $(L+R)$, the $\alpha$ and $\beta$ modes cannot be exactly expressed as a linear superposition of $R$ and $L$. This is because the Möbius strip is a non-orientable manifold on which the chirality of spin precession becomes ambiguous globally, given that $+z$ and $-z$ are indistinguishable. In other words, the eigenspace spanned by $\alpha$ and $\beta$ is not smoothly connected to that spanned by $R$ and $L$ under a global picture; the two cases fall into distinct topological sectors.
The structure of the eigenspace deserves an intuitive explanation. The spin precession on each sublattice can be decomposed into $\ket{\circlearrowright}$ and $\ket{\circlearrowleft}$, while the sublattice degree of freedom is represented by |$\ket{A}$ and |$\ket{B}$. Here $\ket{\circlearrowright}$ ($\ket{\circlearrowleft}$) can be interpreted as $\text{Re}\left[(\hat{x}\pm i\hat{y})e^{i(\omega t-kx)}\right]$. With all constraints relieved, the linear space spanned by the direct products of these vectors is 4D. Under the impact of TBC, however, the magnon eigenspace only covers a 2D subspace. Specifically, the quanta of $\alpha$ and $\beta$ eigenmodes described by Eqs. (17) and (18) can be recast into a suggestive from
@@
where |$\absolutevalue{u_{\rm small}^{\alpha(\beta)}}<\absolutevalue{u_{\rm large}^{\alpha(\beta)}}$ | with $u_{\rm large}^{\alpha(\beta)}+u_{\rm small}^{\alpha(\beta)}\propto\sqrt{q_{\alpha(\beta)}^{1(2)}}$ and $u_{\rm large}^{\alpha(\beta)}-u_{\rm small}^{\alpha(\beta)}\propto\sqrt{q_{\alpha(\beta)}^{2(1)}}$. Comparatively, the eigenspace under the constraint of ordinary PBC, as illustrated in Fig. 3(d), is spanned by